# template.bib

### v1

@@ Abstract
Explaining machine learning (ML) models using eXplainable AI (XAI) techniques has become essential to make them more transparent and trustworthy. This is especially important in high-stakes domains like healthcare, where understanding model decisions is critical to ensure ethical, sound, and trustworthy outcome predictions. However, users are often confused about which explanability method to choose for their specific use case. We present a comparative analysis of widely used explainability methods, Shapley Additive Explanations (SHAP) and Gradient-weighted Class Activation Mapping (GradCAM), within the domain of human activity recognition (HAR) utilizing graph convolutional networks (GCNs). By evaluating these methods on skeleton-based data from two real-world datasets, including a healthcare-critical cerebral palsy (CP) case, this study provides vital insights into both approaches’ strengths, limitations, and differences, offering a roadmap for selecting the most appropriate explanation method based on specific models and applications. We quantitatively and quantitatively compare these methods, focusing on feature importance ranking, interpretability, and model sensitivity through perturbation experiments. While SHAP provides detailed input feature attribution, GradCAM delivers faster, spatially oriented explanations, making both methods complementary depending on the application’s requirements. Given the importance of XAI in enhancing trust and transparency in ML models, particularly in sensitive environments like healthcare, our research demonstrates how SHAP and GradCAM could complement each other to provide more interpretable and actionable model explanations.
Keywords SHAP, GradCAM, XAI, Explainable AI, Human Activity Recognition, HAR
@@ 1 Introduction
Significant progress has been made in the development of Graph Convolution Networks (GCNs) for Human Action Recognition (HAR) in recent years. A significant proportion of the research is dedicated to improving the performance of GCN architectures on widely used benchmark datasets, including NTU RGB+D 60/120 [liuNTURGB+D1202020, shahroudyNTURGB+DLarge2016] and Kinetics [kayKineticsHumanAction2017]. However, less attention has been paid to understanding and explaining the internal workings of the developed architectures. Concurrently, these models’ increasing complexity and scale have reduced their interpretability, turning them into ”black boxes“. Still, understanding the inner workings of such systems is crucial, particularly when they are used in high-risk and sensitive environments such as healthcare, where the outcomes of model decisions can have significant ethical and practical implications [bharatiReviewExplainableArtificial2024, chaddadSurveyExplainableAI2023, lohApplicationExplainableArtificial2022, sadeghiReviewExplainableArtificial2024]. In such contexts, the explainability of models becomes not only a technical concern but also a matter of trust, accountability, and safety [tjoaSurveyExplainableArtificial2021].
@@
In domains like healthcare, where decisions can directly impact patient outcomes, the interpretability of ML models used in this field is critical [koricaExplainableArtificialIntelligence2021, allgaierHowDoesModel2023, nazirSurveyExplainableArtificial2023]. Clinicians require assurance that ML models produce decisions based on relevant, reliable, and ethically sound data patterns [a.SystematicReviewExplainable2023]. The lack of interpretability of such systems could lead to dangerous scenarios, including the perpetuation of biases or the incorrect categorization of patient conditions, which might have significant clinical consequences [gerlingsExplainableAIExplainable2022, kellyKeyChallengesDelivering2019]. Therefore, providing explainability by ”opening the black box“ of GCN-based HAR models is essential for their safe deployment and acceptance in sensitive environments such as healthcare [vellidoImportanceInterpretabilityVisualization2020].
@@
Despite the increasing relevance of XAI, its application within the HAR domain remains largely underexplored. To the extent that XAI methods have been studied for HAR, these focus primarily on three methods: Shapley Additive Explanations (SHAP) [gaoAutomatingGeneralMovements2023, tempelExplainingHumanActivity2024], Class Activation Mapping (CAM) [songStrongerFasterMore2020] and its rectification Gradient-weighted Class Activation Mapping (GradCAM) [dasGradientWeightedClassActivation2022]. Beyond this, there remains a noticeable gap in the literature regarding the systematic comparison of these methods in the HAR field. Specifically, no work has been done to assess which explainability approach is better suited for different types of HAR models and datasets. Further, there is a limited understanding of how different explainability methods vary in their outputs. This gap presents a significant challenge, as the effectiveness and interpretability of XAI methods can vary significantly and depend on the context and model architecture in which they are used. For example, SHAP provides insights into each feature’s importance but may fail to capture the spatial and temporal dynamics in the data. GradCAM, on the other side, can highlight important regions for the model but falls short in explaining the contribution of the individual input features. The lack of a comprehensive comparison of these methods with different datasets and HAR model types poses a challenge in identifying the most appropriate methods for model interpretation. Furthermore, clinicians and researchers should be provided with a quantitative comparison of these two XAI methods to asses which is best suited for their individual HAR model, dataset, and use case. Understanding what to expect from each XAI method regarding explanation quality, transparency, and adaptability is essential to making informed decisions. Hence, a systematic comparison of these two XAI methods is needed. This can pave the way for trustworthy HAR models that can be safely and effectively implemented in real-world applications.
@@ 1.1 Contributions
  • Quantitative comparison of SHAP and GradCAM, in the context of HAR with GCN, on skeleton data from two real-world datasets.
  • Comparative analysis of the two explanation methods, focusing on feature ranking and evaluating the influence of these features through a perturbation sensitivity analysis on predictive performance.
  • A nuanced understanding of the strengths and weaknesses of the two explanation approaches within the HAR domain.
@@ 2 Related Work
2.1 XAI
@@
In the domain of skeleton-based HAR, most XAI techniques rely on saliency-based methods such as CAM [songStrongerFasterMore2020] or GradCAM [dasGradientWeightedClassActivation2022, Song2022ConstructingSA]. Meanwhile, SHAP, which is widely used in a wide range of ML applications, has received less attention in HAR and was only applied in [gaoAutomatingGeneralMovements2023, tempelExplainingHumanActivity2024], to the best of our knowledge.
@@
The first work using XAI within HAR is the work from [songStrongerFasterMore2020]. The authors use CAM to visualize important body key points on the two action classes, throwing and drinking water from the NTU RGB+D dataset [shahroudyNTURGB+DLarge2016]. However, their investigation is restricted to only a qualitative visual representation of the activated body key points for those two action classes. This limits their applicability and quantitative comparability with other techniques. The same authors expand this approach on a refined architecture in [Song2022ConstructingSA], where CAM is used again to produce visual explanations. However, these explanations also lack objective evaluation metrics, making comparing their effectiveness with other approaches difficult. Also, a comparison of the activation maps to the earlier work [songStrongerFasterMore2020] is missing. It can be stated that the CAM approach does not address the contribution of individual input features, an area where SHAP could offer deeper insights. This may be particularly interesting for the architecture used in [Song2022ConstructingSA], where the authors use multiple input branches to process features individually before fusing them into a common stream in the model.
@@
In [gaoAutomatingGeneralMovements2023], SHAP is applied in a medical context for early cerebral palsy (CP) screening. The model combines video data and secondary features like birth weight, sex, and gestational age to detect fidgety movements, indicating an increased risk for neurological dysfunction if they are absent or sporadic [einspielerFidgetyMovementsTiny2016]. These features are then classified and combined into a normal and a risk probability for CP. However, the SHAP explanation is limited to the secondary input features, missing the opportunity to explain the primary input features (i.e., the video data). SHAP would also have to be applied to the primary features to properly explain the model, although this represents a greater technical challenge.
@@
Another work using SHAP within the medical HAR domain is [tempelExplainingHumanActivity2024]. Here, the authors use SHAP to explain the contributions of the primary input features on the model output utilizing two real-world datasets. While this work is the first to apply SHAP on the primary input features, the direct comparison with the often-used GradCAM approach is lacking. A comparison between SHAP and GradCAM could provide a more complete evaluation of the individual strengths and weaknesses. This opens up research on how methods like SHAP and GradCAM might complement or outperform each other in different HAR scenarios and datasets.
@@ 3 Method
This section outlines the two explainability techniques, SHAP and GradCAM, which are compared in this study. Furthermore, the model architecture used to predict the skeleton data is briefly described, including at which layer we obtain the gradients for GradCAM; see a visualization in Fig. 1. Finally, the perturbation technique used to evaluate the model’s sensitivity is explained, including how specific parts of the architecture are perturbed based on the feature ranking from SHAP and GradCAM.
@@
3.1 GradCAM
GradCAM generates an explanation using the gradients from the final convolutional layer of an ML model [selvarajuGradCAMVisualExplanations2020]. To generate explanations, the gradients for class $c$ are computed with respect to the activations of the feature maps in the chosen convolutional layer. Let $y^{c}$ represent the model’s activation for class $c$ before the chosen target layer, corresponding to the model’s prediction for class $c$. The gradients of $y^{c}$ with respect to each feature map activation $A^{k}$, where $A^{k}$ denotes the $k$-th feature map of the final convolutional layer, are then calculated. These gradients indicate how much each spatial location in the feature map contributes to the class prediction. Next, global average pooling is applied to the gradients, yielding the neuron importance weights:
@@ 3.1 GradCAM
The final localization map, which highlights the relevant regions in the skeleton, is then computed as follows:
$L_{\text{GradCAM}}^{c}=\text{ReLU}\left(\sum_{k}\alpha_{k}^{c}A^{k}\right)\,,$ (2)
@@
We additionally apply the counterfactual explanation, where the gradients are negated [selvarajuGradCAMVisualExplanations2020]. This modification allows us to highlight the areas that decrease the class score, identifying the features that are either less important or negatively contributing to, i.e.,  contradict, the prediction. This way, we get the negated importance weights via
@@
and then compute the localization map with (2).
For skeleton-based HAR, GradCAM can be adapted to focus on the relevant spatial-temporal features within the skeletal data. In this context, the skeleton is typically represented as a graph, where nodes correspond to body key points, and edges represent their spatial relations. In GCNs, the final convolutional layer captures the spatial dependencies across the body key points, while the temporal information is encoded across different frames.
In [dasGradientWeightedClassActivation2022], GradCAM is extended to skeleton-based HAR tasks by adapting the method to handle the graph-structured data. Here, the gradients are computed for the graph convolutions, and the localization map indicates which body key points and frames are most influential in the model’s prediction for a specific human activity. This approach can be used to visualize the contributions of different joints and frames, facilitating an understanding of which parts of the body and movement sequences are most relevant for a given activity at a specific time.
@@ 3.2 SHAP
SHAP is a framework for explaining ML model predictions by attributing an importance score to each feature [lundbergUnifiedApproachInterpreting2017]. It is based on the Shapley value from cooperative game theory [shapley7ValueNPerson1997], which measures a player’s contribution to a team’s overall success. In the context of ML, the Shapley value is adapted to evaluate the impact of each input feature on the model’s prediction. The SHAP value $\phi_{i}$ for feature $i$ is computed as
@@
In our study, features are obtained directly from the skeleton data and categorized into four groups: position ($J$), velocity ($V$), bone ($B$), and acceleration ($A$), further described [tempelAutoGCNgenericHumanActivity2024, tempelLightweightNeuralArchitecture2024]. The combined SHAP value of a model trained on these features is thus $\phi=\phi_{P}+\phi_{V}+\phi_{B}+\phi_{A}$. The baseline $\phi_{0}$ is calculated by averaging over $n=100$ randomly sampled data points from the training set $\mathcal{D}_{train}$. Although SHAP is computationally expensive, approximations allow for relatively efficient model explanations. Furthermore, the simple decomposition of the SHAP value into input feature importances provides intuitive explanations that yield a certain interpretability of the model. SHAP values can also be aggregated across multiple instances to provide global insights into feature importances across a dataset, helping to identify critical elements influencing model behavior.

### v2

# xai_shapcam.bib

@@ Abstract
Explaining machine learning (ML) models using eXplainable AI (XAI) techniques has become essential to make them more transparent and trustworthy. This is especially important in high-risk environments like healthcare, where understanding model decisions is critical to ensure ethical, sound, and trustworthy outcome predictions. However, users are often confused about which explanability method to choose for their specific use case. We present a comparative analysis of two explainability methods, Shapley Additive Explanations (SHAP) and Gradient-weighted Class Activation Mapping (Grad-CAM), within the domain of human activity recognition (HAR) utilizing graph convolutional networks (GCNs). By evaluating these methods on skeleton-based input representation from two real-world datasets, including a healthcare-critical cerebral palsy (CP) case, this study provides vital insights into both approaches’ strengths, limitations, and differences, offering a roadmap for selecting the most appropriate explanation method based on specific models and applications. We quantitatively and quantitatively compare the two methods, focusing on feature importance ranking and model sensitivity through perturbation experiments. While SHAP provides detailed input feature attribution, Grad-CAM delivers faster, spatially oriented explanations, making both methods complementary depending on the application’s requirements. Given the importance of XAI in enhancing trust and transparency in ML models, particularly in sensitive environments like healthcare, our research demonstrates how SHAP and Grad-CAM could complement each other to provide model explanations.
Keywords SHAP, Grad-CAM, XAI, Explainable AI, Human Activity Recognition, HAR
@@ 1 Introduction
Significant progress has been made in the development of Graph Convolution Networks (GCNs) for Human Action Recognition (HAR) in recent years. GCNs are considered a practical solution to classify human movement and activities based on skeleton data compared to approaches using raw video data as the input. This is due to the fact that raw video data is often polluted with noise, occluded body parts, and background information that is not informative for action recognition or movement classification. Furthermore, a skeleton representation is free of skin and body composition bias since this information is not embedded in the skeleton sequence. This makes GCNs a practical choice within the HAR domain if the skeleton representation is available. A considerable proportion of the research is dedicated to improving the performance of GCN architectures on widely used benchmark datasets, including NTU RGB+D 60/120 [liuNTURGB+D1202020, shahroudyNTURGB+DLarge2016] and Kinetics [kayKineticsHumanAction2017]. However, less attention has been paid to understanding and explaining the internal workings of the developed architectures. Concurrently, these models’ increasing complexity and size have reduced the explainability aspect, turning them into “black boxes”. Still, understanding the inner workings of such systems is crucial, particularly when they are used in high-risk and sensitive environments such as healthcare, where the outcomes of model decisions can have significant ethical and practical implications [bharatiReviewExplainableArtificial2024, chaddadSurveyExplainableAI2023, lohApplicationExplainableArtificial2022, sadeghiReviewExplainableArtificial2024]. In such contexts, the explainability of models becomes not only a technical concern but also a matter of trust, accountability, and safety [tjoaSurveyExplainableArtificial2021].
@@
In domains like healthcare, where decisions can directly impact patient outcomes, the explainability of ML models used in this field is critical [koricaExplainableArtificialIntelligence2021, allgaierHowDoesModel2023, nazirSurveyExplainableArtificial2023]. Clinicians require assurance that ML models produce decisions based on relevant, reliable, and ethically sound data patterns [a.SystematicReviewExplainable2023]. The lack of explainability of such systems could lead to dangerous scenarios, including the perpetuation of biases or the incorrect categorization of patient conditions, which might have significant clinical consequences [gerlingsExplainableAIExplainable2022, kellyKeyChallengesDelivering2019]. Therefore, providing explainability by “opening the black box” of GCN-based HAR models is essential for their safe deployment and acceptance in sensitive environments such as healthcare [vellidoImportanceInterpretabilityVisualization2020].
@@
Despite the increasing relevance of XAI, its application within the HAR domain remains largely underexplored. To the extent that XAI methods have been studied for HAR, these focus primarily on three methods: Shapley Additive Explanations (SHAP) [gaoAutomatingGeneralMovements2023, tempelExplainingHumanActivity2025], Class Activation Mapping (CAM) [songStrongerFasterMore2020] and its rectification Gradient-weighted Class Activation Mapping (Grad-CAM) [dasGradientWeightedClassActivation2022]. Beyond this, there remains a noticeable gap in the literature regarding the systematic comparison of these methods in the HAR field. Specifically, no work has been done to assess which explainability approach is better suited for different types of HAR models and datasets. Further, there is a limited understanding of how different explainability methods vary in their outputs. This gap presents a significant challenge, as the effectiveness and explainability of XAI methods can vary significantly and depend on the context and model architecture in which they are used. For example, SHAP provides insights into each feature’s importance but may fail to capture the spatial and temporal dynamics in the data. Grad-CAM can highlight important regions for the model but may fall short of explaining the contribution of the individual input features. The lack of a comprehensive comparison of these methods with different datasets and HAR model types poses a challenge in identifying the most appropriate methods for model explanation. Furthermore, clinicians and researchers should be provided with a quantitative comparison of these two XAI methods to assess which is best suited for their individual HAR model, dataset, and use case. Understanding what to expect from each XAI method regarding explanation quality, transparency, and adaptability is essential to making informed decisions. Hence, a systematic comparison of these two XAI methods is needed. This can pave the way for trustworthy HAR models that can be safely and effectively implemented in real-world applications.
@@ 1.1 Contributions
  • Expanding an existing XAI framework to provide Grad-CAM explanations.
  • First work to compare two explanation methods, SHAP and Grad-CAM, within the HAR domain, focusing on feature ranking, quantitative evaluation of the explanation’s reliability, and a qualitative comparison.
  • A nuanced understanding of the strengths and weaknesses of the two explanation approaches for GCNs within the HAR domain.
@@ 2 Related Work
2.1 XAI Methods
To provide explanations of ML models, a range of XAI methods have been applied [bharatiReviewExplainableArtificial2024, chaddadSurveyExplainableAI2023, tjoaSurveyExplainableArtificial2021, allgaierHowDoesModel2023]. Among the most commonly used ones within the HAR domain are SHAP and Grad-CAM [lundbergUnifiedApproachInterpreting2017, selvarajuGradCAMVisualExplanations2020], which we apply in this study and discuss in greater detail in Sec. 3.1 and 3.2. Other approaches within XAI include saliency maps [simonyanDeepConvolutionalNetworks2014] or Local Interpretable Model-agnostic Explanations (LIME) [ribeiroWhyShouldTrust2016].
Saliency maps highlight the regions of an input that most influence a model’s decision [simonyanDeepConvolutionalNetworks2014]. These visualizations help identify key input features and can guide dimensionality reduction, as shown in [yanResNetLikeCNNArchitecture2022]. While the use of saliency maps has been well validated in computer vision tasks [simonyanDeepConvolutionalNetworks2014, adebayoSanityChecksSaliency2018], direct applications in skeleton-based HAR are sparse, to the best of our knowledge. Furthermore, saliency maps often provide a broad, unstable view of the model’s decision-making process [adebayoSanityChecksSaliency2018].
LIME uses interpretable surrogate models to explain predictions [ribeiroWhyShouldTrust2016]. It does this by generating small input variations and training a simplified model to approximate how the complete model behaves locally [ribeiroWhyShouldTrust2016]. While LIME has proven effective across various domains, its application in HAR remains unexplored, to the best of our knowledge. The challenge in applying LIME within HAR lies in generating kinematic sound perturbations for skeleton-based data since random changes to the body coordinates can easily create unrealistic human movements, making the explanations unreliable.
In summary, while various XAI methods offer promising tools for interpreting ML models, their adaptation to skeleton-based HAR remains an open research challenge, especially when preserving kinematic sound data is necessary.
2.2 XAI in Human Activity Recognition
@@
In the domain of skeleton-based HAR, most XAI techniques rely on saliency-based methods such as CAM [songStrongerFasterMore2020] or Grad-CAM [dasGradientWeightedClassActivation2022, Song2022ConstructingSA]. Meanwhile, SHAP, which is widely used in a wide range of ML applications, has received less attention in HAR and was only applied in [gaoAutomatingGeneralMovements2023, tempelExplainingHumanActivity2025], to the best of our knowledge.
@@
The first work using XAI within HAR is the work from [songStrongerFasterMore2020]. The authors use CAM to visualize important body key points on the two action classes, throwing and drinking water from the NTU RGB+D dataset [shahroudyNTURGB+DLarge2016]. However, their investigation is restricted to only a qualitative visual representation of the activated body key points for those two action classes. This limits their applicability and quantitative comparability with other techniques. The same authors expand this approach on a refined architecture in [Song2022ConstructingSA], where CAM is used again to produce visual explanations. However, these explanations also lack objective evaluation metrics, making comparing their effectiveness with other approaches difficult. Also, a comparison of the activation maps to the earlier work [songStrongerFasterMore2020] is missing. It can be stated that the CAM approach does not address the contribution of individual input features, an area where SHAP could offer more profound insights. This may be particularly interesting for the architecture used in [Song2022ConstructingSA], where the authors use multiple input branches to process features individually before fusing them into a common stream in the model.
@@
In [gaoAutomatingGeneralMovements2023], SHAP is applied in a medical context for early cerebral palsy (CP) screening. The model combines video data and secondary features like birth weight, sex, and gestational age to detect fidgety movements, indicating an increased risk for neurological dysfunction if they are absent or sporadic [einspielerFidgetyMovementsTiny2016]. These features are then classified and combined into a normal and a risk probability for CP. However, the SHAP explanation is limited to the secondary input features, missing the opportunity to explain the primary input features (i.e., the video data). SHAP would also have to be applied to the primary features to adequately explain the model, although this represents a more significant technical challenge.
@@
Another work using SHAP within the medical HAR domain is [tempelExplainingHumanActivity2025], which this work builds upon. Here, the authors use SHAP to explain the contributions of the primary input features on the model output utilizing two real-world datasets. While this work is the first to apply SHAP on the primary input features, the direct comparison with the often-used Grad-CAM method within HAR is lacking. A comparison between SHAP and Grad-CAM can provide a more complete evaluation of the individual strengths and weaknesses, which is not possible if only one explanation method is used. This opens up the question of how methods like SHAP and Grad-CAM might complement or outperform each other in different HAR scenarios and datasets.
@@ 3 Method
This section outlines the two explainability techniques, SHAP and Grad-CAM, which are compared in this study. Furthermore, the model architecture used to predict the skeleton data is briefly described, including at which layer we obtain the gradients for Grad-CAM; see a visualization in Fig. 1. Finally, the perturbation technique used to evaluate the model’s sensitivity is explained, including how specific parts of the architecture are perturbed based on the feature ranking from SHAP and Grad-CAM.
@@
3.1 Grad-CAM
Grad-CAM generates an explanation using the gradients from the final convolutional layer of an ML model [selvarajuGradCAMVisualExplanations2020]. This is done by computing the gradients for class $c$ with respect to the activations of the feature maps in the chosen convolutional layer. Let $y^{c}$ represent the model’s activation for class $c$ before the chosen target layer, corresponding to the model’s prediction for class $c$. The gradients of $y^{c}$ are calculated concerning each feature map activation $A^{k}$, where $A^{k}$ denotes the $k$-th feature map of the final convolutional layer. These gradients indicate how much each spatial location in the feature map contributes to the class prediction. Next, global average pooling is applied to the gradients, yielding the neuron importance weights
@@ 3.1 Grad-CAM
The final localization map, which highlights the relevant regions in the skeleton, is then computed as
$L_{\text{Grad-CAM}}^{c}=\text{ReLU}\left(\sum_{k}\alpha_{k}^{c}A^{k}\right)\,,$ (2)
@@
We additionally apply the counterfactual explanation, where the gradients are negated [selvarajuGradCAMVisualExplanations2020]. This modification allows us to highlight the areas that decrease the class score, identifying the features that are either less important or negatively contribute to the prediction, i.e.,  contradict. This way, we get the negated importance weights via
@@
and then compute the localization map with eq. (2).
For skeleton-based HAR, Grad-CAM can be adapted to focus on the relevant spatial-temporal features within the skeletal data. In this context, the skeleton is typically represented as a graph, where nodes correspond to body key points, and edges represent their spatial relations. In GCNs, the final convolutional layer captures the spatial dependencies across the body key points, while the temporal information is encoded across different frames.
In [dasGradientWeightedClassActivation2022], Grad-CAM is extended to skeleton-based HAR tasks by adapting the method to handle the graph-structured data. Here, the gradients are computed for the graph convolutions, and the localization map indicates which body key points and frames are most influential in the model’s prediction for a specific human activity. This approach can be used to visualize the contributions of different joints and frames, facilitating an understanding of which parts of the body and movement sequences are most relevant for a given activity at a specific time.
@@ 3.2 SHAP
SHAP is a framework for explaining ML model predictions by attributing an importance score to each feature [lundbergUnifiedApproachInterpreting2017]. It is based on the Shapley value from cooperative game theory [shapley7ValueNPerson1953], which measures a player’s contribution to a team’s overall success. In the context of ML, the Shapley value is adapted to evaluate the impact of each input feature on the model’s prediction. The SHAP value $\phi_{i}$ for feature $i$ is computed as
@@
In our study, features are obtained directly from the skeleton data and categorized into four groups: position ($J$), velocity ($V$), bone ($B$), and acceleration ($A$), further described [tempelAutoGCNgenericHumanActivity2024, tempelLightweightNeuralArchitecture2024]. The combined SHAP value of a model trained on these features is thus $\phi=\phi_{J}+\phi_{V}+\phi_{B}+\phi_{A}$. The baseline $\phi_{0}$ is calculated by averaging over $n=100$ randomly sampled data points from the training set $\mathcal{D}_{train}$. Although SHAP is computationally expensive, approximations allow for relatively efficient model explanations. Furthermore, the simple decomposition of the SHAP value into input feature importances provides intuitive explanations of the model. SHAP values can also be aggregated across multiple instances to provide global insights into feature importances across a dataset, helping to identify critical elements influencing model behavior.